Use the binomial formula to find the coefficient of the z^y term in the expansion of (2z+y). E $ ?
Answer #1

Hey there stay home and stay safe

\small \\$We need to find the coefficients of $z^4y^9 $ term in the expansion of $ (2z+y)^{13}\\ (a+b)^n=\ ^nC_0a^nb^0+^nC_1a^{n-1}b^1+....+^nC_{p}a^{n-p}b^p+...+^nC_na^0b^n\\ $The power of the first term z is 4\\ The term consisting $z^4 $ is $ ^{13}C_9(2z)^4y^9=^{13}C_92^4z^4y^9\\ $So the coefficient is $^{13}C_92^4=11440

And please leave a like. Thank you.